State, giving reasons, whether the following statements are True or False:
(1) When a dielectric material is inserted between the parallel plates of a capacitor, the
electrostatic energy stored between the plates increases.
1
Expert's answer
2020-02-17T09:20:46-0500

If a plate having the plate area A of the parallel plate capacitor,

d is the distance between the plates.

So, Capacitance

"C_o=\dfrac{\epsilon_o A}{d}"

when the dielectric material is inserted between the plates.

Let the di electric constant of the material =K

"C=\dfrac{K\epsilon_o A}{d}"

"C=KC_o"

We know that the energy stored in the capacitor is

"E=\dfrac{K\epsilon AV^2}{2d}"

"E\propto kC_o"

So, it can store more charges and maintain much potential difference.

Hence energy of the plates will increase.